Telescope and AGw Commissioning at LFBG with IRTC - 29-30 April 2009 MST

Observers:: R .Green, J. Heidt
Telescope Operator: David Gonzalez Huerta
Software Support: (phone) Telescope Support:

Summary

  • Good night overall as far as weather is concerned. Seeing mostly ~0.5-0.8" on guider in r'.
  • Severe problems with encoders, see Issue track#2068, lowered observing efficiency substantially
  • Offset tests with Daves dice5-script, 2"
  • Data for new pointing model

Preparations

JMH has installed a new SX Primary Mirror Collimation Table based on last night's data. Press Initialize on the PSF Primary GUI to load it. See the discussion and plots at the bottom of last night's log - Com20090429.

Details

Open

19:00 MST Dome opened

19:30 Tdome=5.8 T(M1)=3.2 Tout=5.7 RH=22%, wind=10m/s

19:48 pointto M5_0347 Tel="TRACK" PA=0 PAM-.

PTspiral, no star found

20:40 pointto M5_0425 Tel="ACQUIRE" gs=0 PA=0 PAM- (to see images on the GC)
PTspiral, OK, star found, CA=-12 IE=-93

20:34 pointto M5_0384 Tel="ACQUIRE" gs=0 PA=0 PAM- PTspiral, still no star found
Reset AZ.encoders

21:06 pointto M5_0384 Tel="ACQUIRE" gs=0 PA=0 PAM- PTspiral, again no success

21:13 pointto Regulus PTspiral, again no success
Reset EL.encoders

21:34 Pointto Regulus again => successfull

21: 35 pointto M5_0384 again => YES, CA=-69, IE=-90

21:42 pointto BS9148
TO work station got sick, reboot TCS + start networks,

21:54 pointto BS9148 again for initial collimation, far out of focus
WFS-RMS: 2820, 800, 2950, 1750, 257, 180

22:09 pointto BS9148 gs=5, lost guide star

22:15 M5_1332, recheck pointing, CA=-60, IE=-64

22:09 pointto BS9148 gs=5, OK

22:29 takepic BS9148, 052945 to see that everything works
22:29 takeseq BS9148 00001_cube to see that everything works

Offset-tests (small offsets)

22:32 dice5 offsets 2",2" (DETXY), 1, 45 loop=5
053436, cubes 2-5
stopped after 4 cubes, problems with guiding...

22:39 Preset, and started again,
053944-055634, cubes 6-18, only 13 cubes saved???
Received several times a warning: "GCS server: ERROR: failed to move to offset position"
"GUIDING OK"
Guiding camera reports seeing around 0.5"
Moon still up, very few cirrus around

22:59 Preset again, with gs=5, but PA=indef (was set=0 before :-=(
060850-, cubes, stopped, gs=5 bad choice, only 0.3" from center
Everything in IRAF got stuck => restart

23:19 preset BS9148, gs=4, tried dice5 (DETXY) again, seems to work now
062306-065542, cubes 25-48, OK

23:58 preset BS9148, gs=4, dice 5 (RADEC)
065924-073126, cubes 49-72

Pointing model

00:40 preset BS9153, determine new rotator center
PA=0, 074338, cube 73
PA=180, 074522, cube 74

00:45 Update collimation, OK

00:52 Filename pointing log: 20090430.ptlog

00:55 Update pointing, CA=-59, IE=-74

01:06 Start pointing model observations:

01 ACT0356, 080930, cube 75
ACT0405, 081216, cube 76, star not there, do PTspiral
02 ACT0405, 083329, cube 77, CA=-63, IE=-125
03 WT10_399, 083836, cube 78, CA=-73, IE=-140
04 WT10_275, 084205, cube 79
05 ACT0171, 084532, cube 80
06 WT10_245, 085008, cube 81, Ca=-73 IE=-150
07 ACT0341, 085714, cube=82, CA=-78, IE=-140
WT10_077, (EL=18degs, dec=-38), not found via PTspiral, skipped
Switched from strip.enconder to motor.encoder and back,
Do that after future presets in case of large errors (mostly the case)
=> first 7 measurements probably not useful since decoder errors large
Doing new collimation on BS9178
08 WT10_179, 092926, cube 83
09 WT10_077, 093601, cube 84, tried again, found it, CA=-60, IE=-70
10 WT10_268, 094122, cube 85, CA=-69, IE=-68
11 WT10_219, 094538, cube 86, IE=-80, CA=-60
12 WT10_321, 094938, cube 87, IE=-86, CA=-68
13 ACT0334, 095212, cube 88
14 WT10_304, 095625, cube 89, IE=-90, CA=-68
Doing new collimation on BS9166
15 WT10_300, 101436, cube 90, IE=-87, CA=-60
16 WT10_272, 101855, cube 91, IE=-90, CA=-69
17 ACT0179, 102312, cube92, IE=-80, AC=-60
18 WT10_094, 102620, cube 93, IE=-70, CA=-60
19 WT10_109, 102850, cube 94
20 ACT0300, 103353, cube 95, IE=-85, CA=-50
21 WT10_264, 103803, cube 96, IE=-70, CA=-69
22 WT10_207, 104319, cube 97, IE=-65, CA=-50
23 WT10_311, 104727, cube 98, IE=-60, CA=-75
24 ACT0330, 105104, cube 99
25 ACT0396, 110856, cube 100, IE=-80, CA=-69
26 ACT0404, 111319, cube 101, IE=-76, CA=-60
Doing new collimation on BS9177, IE=-80, CA=-69
27 ACT0411, 112751, cube 102 (GProbe not out), 112931, cube 103 (ok)
28 ACT0332, 113319, cube 104
29 ACT0382, 113642, cube 105
30 WT10_386, 114100, cube 106, IE=-75, CA=-63
31 WT10_405, 114604, cube 107, IE=-79, CA=-80
32 WT10_411, 114914, cube 108
33 WT10_369, 115228, cube 109
34 WT10_385, 115759, cube 110, IE=-80, CA=-65
35 ACT0354, 120112, cube 111
36 ACT0356, 120424, cube 112
37 ACT0315, 120828, cube 113 , IE=-65, CA=-80
38 WT10_300, 121103, cube 114
39 WT10_211, 121511, cube 115, IE=-75, CA=-70

07:30 MST 30 April 2009 - pointing model generated. The first 7 points did indeed need to be deleted. The remaining stars had 5.5 arcsec rms on the fit with only the standard ALTAZ model plus tube flexure and a harmonic in AZ proportional to cos(2*AZ). The AZ scale error was consistent with zero.

Close

05:15 Done, closing dome.

-- JohnHill/Jochen Heidt - 30 Apr 2009
Topic revision: r8 - 01 May 2009, JochenHeidt
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback